# Evidence for Intrinsic Galaxy Alignments in Ellipticity Autocorrelations out to $100\,h^{-1}\,{\rm Mpc}$ from SDSS Galaxies with DESI Imaging

### v1

@@ Abstract
Measuring the autocorrelation of galaxy shapes, the so-called intrinsic-intrinsic (II) correlation is important for both cosmology and understanding the formation of massive elliptical galaxies. However, such measurements are significantly more challenging than the cross-correlation with galaxy density (GI correlation) due to the much lower signal-to-noise ratio. In this paper, we report the first observational evidence for large-scale intrinsic alignments measured from the ellipticity autocorrelations, extending out to $100\,h^{-1}\,{\rm Mpc}$. From SDSS and SDSS-III BOSS, we analyze over the redshift range $0.16\leq z\leq 0.70$ luminous red galaxies (LRG), LOWZ and CMASS galaxy samples, the latter two of which are cross-matched with high-quality DESI imaging data. By expanding one of the two II correlation functions, $\mathrm{II}(-)$, in terms of the associated Legendre polynomials, we effectively isolate the line-of-sight projection effects and enhance the signal. The resulting correlation for all three samples exhibits a clear power-law form. We also show that combining the two II correlations, $\mathrm{II}(+)$ and $\mathrm{II}(-)$, increases the detection significance by $\sim 10\%$, though they appear equivalent mathematically under the linear alignment model. Importantly, this measurement opens a new observational window for probing signals uniquely encoded in shape autocorrelations, such as tensor perturbations from the gravitational waves. This analysis establishes a practical framework for probing such effects.
Observational cosmology — Large-scale structure of universe — Redshift surveys — Gravitational lensing
@@ 1 Introduction
Intrinsic alignments (IA) of galaxy shapes are a key phenomenon linking large-scale structure with galaxy formation and offering a complementary probe of cosmology. Originally proposed as a contaminant to weak gravitational lensing measurements, the autocorrelation of galaxy shapes, known as the intrinsic–intrinsic shape (II) correlation, was first studied (Heavens et al., 2000; Croft & Metzler, 2000; Lee & Pen, 2000; Catelan et al., 2001; Jing, 2002; Crittenden et al., 2002). Hirata & Seljak (2004) later emphasized that the cross-correlation between galaxy positions and intrinsic shape (GI correlation) can be an even more serious systematic. Because GI correlations are easier to measure observationally, especially using bright galaxies or clusters, they have become the dominant focus in large-scale IA studies, both theoretically and observationally (Mandelbaum et al., 2006; Hirata et al., 2007; Okumura & Jing, 2009; Joachimi et al., 2011; Smargon et al., 2012; Li et al., 2013; Singh et al., 2015; van Uitert & Joachimi, 2017; Tonegawa & Okumura, 2022; Desai & Ryden, 2022; Xu et al., 2023a; Hervas Peters et al., 2024; Shi et al., 2024; Lamman et al., 2024). In contrast, II correlations, which directly probe alignments between galaxy shapes, are significantly noisier and more difficult to detect at large scales. Consequently, only a few studies have succeeded in measuring them with high significance. Okumura et al. (2009) reported the clearest detection to date, finding an ellipticity autocorrelation signal out to $\sim 30\,h^{-1}\,{\rm Mpc}$ using luminous red galaxies (LRGs) from the Sloan Digital Sky Survey (SDSS).
@@
Recently, the potential of IA measurements as cosmological probes has attracted growing attention. In particular, the use of ellipticity correlations in three dimensions has been shown to enable constraints on redshift-space distortions (RSD) and baryon acoustic oscillations (BAO), offering both dynamical and geometric information (Chisari & Dvorkin, 2013; Okumura et al., 2019; Taruya & Okumura, 2020; Okumura & Taruya, 2022; Chuang et al., 2022; Lee et al., 2023; Saga et al., 2023; Shim et al., 2025). However, observational constraints at large scales continue to rely almost exclusively on GI correlations (Okumura & Taruya, 2023; Kurita & Takada, 2023; Xu et al., 2023b), due to the difficulty of extracting clean II signals from data.
@@
Despite the challenges, measuring shape autocorrelations remains essential because they carry unique information inaccessible to GI cross-correlations. Particularly, shape autocorrelations play a critical role in probing new physics, such as the stochastic gravitational wave (GW) background, through their imprint on galaxy orientations (Schmidt et al., 2014; Biagetti & Orlando, 2020; Akitsu et al., 2023a; Philcox et al., 2024; Saga et al., 2024; Okumura & Sasaki, 2024). Recent work showed that IA signals projected along the line of sight are more naturally expanded using the associated Legendre basis rather than the standard Legendre one, due to the geometric anisotropy of projected shapes (Kurita & Takada, 2022; Okumura et al., 2024; Singh et al., 2024; Inoue et al., 2025). This formalism improves the separation of physical signals from noise and enhances interpretability at large scales.
In this paper, we measure the II correlation functions using various galaxy samples from the SDSS and SDSS-III Baryon Oscillation Spectroscopic Survey (BOSS), covering the redshift range $0.16\leq z\leq 0.70$. To maximize the signal-to-noise ratio (S/N), we expand the II correlations in the associated Legendre basis. Following Xu et al. (2023b), we further improve the shape measurements by cross-matching the BOSS spectroscopic galaxy samples with Dark Energy Spectrograph Instrument (DESI) imaging data. This allows us to report the first observational evidence for intrinsic galaxy alignments in ellipticity autocorrelations out to $100\,h^{-1}\,{\rm Mpc}$, a scale comparable to that of the BAO. Our results demonstrate that galaxy shape autocorrelations can be robustly measured at cosmological scales, opening a new window for using IA as a cosmological probe. Throughout this paper, we adopt a flat $\Lambda$CDM model determined by Planck Collaboration et al. (2020).
@@ 2 SDSS and DESI Galaxy Samples
where $q$ is the minor-to-major-axis ratio ($0\leq q\leq 1$) and $\beta$ is the position angle of the ellipticity from the north celestial pole to the east. To improve the shape measurement quality for the LOWZ and CMASS galaxies, we follow the method of Xu et al. (2023b) and cross-match these samples with the DESI Legacy Imaging Survey DR9 data (DESI Collaboration et al., 2016; Dey et al., 2019). The DESI imaging covers the entire BOSS footprint and provides imaging data two to three magnitudes deeper than the SDSS photometry used in target selection, sufficient to robustly determine galaxy orientations. For the LRG sample, we use the original shape measurements defined by the $25~{}{\rm mag}~{}{\rm arcsec}^{-2}$ isophote in the $r$ band, which has been shown to be effective for IA analysis. Accordingly, we adopt the same LRG dataset as in Okumura & Taruya (2023).
The total numbers of the LRG, LOWZ, and CMASS used in this study are $105,334$, 355781, and 737435, respectively. As in our previous IA analyses, we set the axis ratio to $q=0$ in Equation (5) (Okumura & Jing, 2009; Okumura et al., 2009, 2019, 2020). Since our primary interest is in the statistical accuracy of the shape correlations, rather than the amplitude of IA, this simplification does not affect our conclusions.
Figure 1: Multipole moments of the intrinsic ellipticity autocorrelation functions expanded in the standard Legendre basis, measured from SDSS LRG (left), LOWZ (middle), and CMASS (right) galaxy samples. The upper set shows the $\mathrm{II}(+)$ correlation multipoles, $\xi_{+,\ell}$, while the lower set shows the $\mathrm{II}(-)$ multipoles, $\xi_{-,\ell}$, for $\ell=0$ (monopole; circles) and $\ell=2$ (quadrupole; triangles). The correlation amplitude for the LRG sample is multiplied by $1/4$. The LRG results are reproduced from Okumura & Taruya (2023), while the LOWZ and CMASS measurements incorporate high-quality DESI imaging data, significantly improving the signal. The solid curves are the best-fit nonlinear model fitted using the data points enclosed by the vertical lines. The dotted curves are the linear predictions.
@@ 3 Measurement of Correlation Functions
$\xi_{\pm}({\boldsymbol{r}})=\frac{S_{+}S_{+}\pm S_{\times}S_{\times}}{RR}\ ,$ (7)
where $RR$ is the normalized count of random-random pairs, and $S_{+}S_{+}=\sum_{i\neq j|{\boldsymbol{r}}}{\gamma_{+}(j|i)\gamma_{+}(i|j)}$ and similarly for $S_{\times}S_{\times}$.
Since the IA correlations are anisotropic functions, we bin galaxy pairs in both radial separation $r=|\boldsymbol{r}|$ and the cosine of the angle to the line of sight, $\mu_{\boldsymbol{r}}$. We then compute the multipole moments of the correlation functions using the Legendre expansion:
@@
where ${\cal L}_{\ell}$ is the $\ell$th-order Legendre polynomial and $\Delta\mu_{\boldsymbol{r}}=0.1$ is the angular bin size. Due to the geometric factor arising from the projection of the shape field along the line of sight (see Eq. (16) below), the $\mathrm{II}(-)$ correlation is more naturally expanded in terms of the associated Legendre polynomials,
@@
where $\Theta_{L}^{m}(\mu)$ is the normalized associated Legendre function related to the unnormalized one by $\Theta_{L}^{m}(\mu)=\sqrt{\frac{2L+1}{2}\frac{(L-m)!}{(L+m)!}}\mathcal{L}_{L}^{m}(\mu_{\boldsymbol{k}})$ and $L\geq m=4$. We use a tilde to distinguish these coefficients from those expanded in the standard Legendre basis. The summation in Eqs. (8) and (9) is performed over $-1\leq\mu_{\boldsymbol{r}}\leq 1$.
Figure 2: Upper panels: Similarly to Fig. 1, but $L=4$ moments of $\mathrm{II}(-)$ correlation functions of SDSS galaxies expanded by the associated Legendre basis, $\widetilde{\xi}_{-,L}$. The solid curves are the best-fit nonlinear alignment and RSD models fitted using the data points enclosed by the vertical lines. Lower panels: Same as upper panels but the vertical axes are multiplied by $r^{2}$ so that they can be directly compared to the statistics shown in Fig. 1.
@@
In this analysis, we focus only on the two lowest multipoles: the monopole ($\ell=0$) and quadrupole ($\ell=2$), as the hexadecapole ($\ell=4$) is too noisy. Even for the lower multipoles, it is non-trivial to argue up to which scale the correlation is detected. The monopoles of $\mathrm{II}(+)$ appear non-zero for all three samples up to several tens of $\,h^{-1}\,{\rm Mpc}$, while all the quadrupoles remain statistically consistent with zero. For the $\mathrm{II}(-)$ correlation, the monopoles are better measured than those of $\mathrm{II}(+)$. The detection for LRG up to $\sim 40\,h^{-1}\,{\rm Mpc}$ corresponds to the earlier finding of Okumura et al. (2009). Notably, the LOWZ and CMASS results are significantly improved compared to those in Okumura & Taruya (2023), due to the enhanced imaging depth and quality provided by the DESI data. The quadrupole of $\mathrm{II}(-)$ for CMASS appears non-zero over a relatively large range of scales. However, because the signal is distributed over multiple multipoles and the overall measurements remain noisy, these results do not allow a definitive detection based solely on the standard Legendre expansion.
@@ 4.1 Nonlinear Alignment and RSD Model
Table 1: Constraints on the amplitude of IA and $\chi^{2}$ values.
@@
Statistics dof $-b_{K}\sigma_{8}$ $\chi_{\min}^{2}/{\rm dof}$ $-b_{K}\sigma_{8}$ $\chi_{\min}^{2}/{\rm dof}$ $-b_{K}\sigma_{8}$ $\chi_{\min}^{2}/{\rm dof}$
@@
where $\delta_{m}$ is the nonlinear matter density field and $b_{K}$ is the redshift-dependent coefficient of the IA, which we refer to as the shape bias. Furthermore, the redshift-space shape field is multiplied by the damping function due to the Finger-of-God (FoG) effect. We then define E-/B-modes, $\gamma_{(E,B)}$, which are the rotation-invariant decomposition of the ellipticity field (Crittenden et al., 2002),
@@
where $k=|\boldsymbol{k}|$, $\mu_{\boldsymbol{k}}$ is the direction cosine between the line of sight and the wavevector $\boldsymbol{k}$, and $P_{\delta\delta}$ is the nonlinear matter power spectrum computed by the revised Halofit (Takahashi et al., 2012).
Since the FoG effect comes into IA power spectra in the same way as the conventional galaxy power spectrum, we adopt a simple Gaussian function to characterize the FoG damping, $D_{\rm FoG}(k\mu_{\boldsymbol{k}}\sigma_{v})=\exp{\left(-k^{2}\mu_{\boldsymbol{k}}^{2}\sigma_{v}^{2}/2\right)}$, with $\sigma_{v}$ the nonlinear velocity dispersion parameter. Since $P_{\delta\delta}$ is proportional to the square of the normalization parameter of the density fluctuation, $\sigma_{8}^{2}(z)$, free parameters for this model are ${\boldsymbol{\theta}}=(b_{K}\sigma_{8},\sigma_{v})$.
To take into account the line-of-sight anisotropy in Eq. (16), we use the multipole expansion in terms of the standard or associated Legendre polynomials (Kurita & Takada, 2022; Okumura et al., 2024),
@@
$\xi_{+,\ell}(r)=\mathcal{H}_{\ell}^{-1}\left[P_{EE,\ell}(k)\right](r)~{},$ (19)
$\widetilde{\xi}_{-,L}(r)=\mathcal{H}_{L}^{-1}\left[\widetilde{P}_{EE,L}(k)\right](r)~{},$ (20)
$\mathcal{H}_{\ell}^{-1}\left[g(k)\right](r)\equiv i^{\ell}\int\frac{k^{2}dk}{2\pi^{2}}j_{\ell}(kr)g(k)~{}.$ (21)
Converting the $\mathrm{II}(-)$ correlation function from the associated to the standard Legendre bases can be performed via the linear transform,
@@
where the coefficients $A_{\ell L}$ are constants and provided in Appendix A. In the linear theory limit, $\sigma_{v}\to 0$ and thus $D_{\rm FoG}=1$. Hence, Equation (16) converges to the linear theory formulas derived in Okumura & Taruya (2020) (see also Appendix C of Okumura et al., 2024).

### v2

@@ Abstract
Measuring the autocorrelation of galaxy shapes, known as the intrinsic-intrinsic (II) correlation, is important for both cosmology and understanding the formation of massive elliptical galaxies. However, such measurements are significantly more challenging than those of the cross-correlation with galaxy density (GI correlation) due to the much lower signal-to-noise ratio. In this Letter, we report the first observational evidence for large-scale intrinsic alignments measured from the ellipticity autocorrelations, extending out to $100\,h^{-1}\,{\rm Mpc}$. From the Sloan Digital Sky Survey (SDSS) and SDSS-III Baryon Oscillation Spectroscopic Survey, we analyze, over the redshift range $0.16\leq z\leq 0.70$, luminous red galaxy, LOWZ, and CMASS galaxy samples, the latter two of which are crossmatched with high-quality Dark Energy Spectrograph Instrument imaging data. By expanding one of the two II correlation functions, $\mathrm{II}(-)$, in terms of the associated Legendre polynomials, we effectively isolate the line-of-sight projection effects and enhance the signal. The resulting correlation for all three samples exhibits a clear power-law form. We also show that jointly analyzing the two II correlations, $\mathrm{II}(+)$ and $\mathrm{II}(-)$, increases the detection significance by $\sim 10\%$, even though both are derived from the same $E$-mode power spectrum. Importantly, this measurement opens a new observational window for probing signals uniquely encoded in shape autocorrelations, such as tensor perturbations from the gravitational waves. Our analysis establishes a practical framework for extracting such effects.
[…]
journal: ApJ Letters
[…]
August 28, 2025
show]tokumura@asiaa.sinica.edu.tw
@@ 1 Introduction
Intrinsic alignments (IAs) of galaxy shapes are a key phenomenon linking large-scale structure with galaxy formation and offering a complementary probe of cosmology. Originally proposed as a contaminant to weak gravitational lensing measurements, the autocorrelation of galaxy shapes, known as the intrinsic–intrinsic shape (II) correlation, was first studied (Croft & Metzler, 2000; Heavens et al., 2000; Lee & Pen, 2000; Catelan et al., 2001; Crittenden et al., 2002; Jing, 2002). Hirata & Seljak (2004) later emphasized that the cross correlation between galaxy positions and intrinsic shape (GI correlation) can be an even more serious systematic. Because GI correlations are easier to measure observationally, especially using bright galaxies or clusters, they have become the dominant focus in large-scale IA studies, both theoretically and observationally (Mandelbaum et al., 2006; Hirata et al., 2007; Okumura & Jing, 2009; Joachimi et al., 2011; Smargon et al., 2012; Li et al., 2013; Singh et al., 2015; van Uitert & Joachimi, 2017; Desai & Ryden, 2022; Tonegawa & Okumura, 2022; Xu et al., 2023a; Lamman et al., 2024; Shi et al., 2024; Hervas Peters et al., 2024). In contrast, II correlations, which directly probe alignments between galaxy shapes, are significantly noisier and more difficult to detect at large scales. Consequently, only a few studies have succeeded in measuring them with high significance. Okumura et al. (2009) reported the clearest detection to date, finding an ellipticity autocorrelation signal out to $\sim 30\,h^{-1}\,{\rm Mpc}$ using luminous red galaxies (LRGs) from the Sloan Digital Sky Survey (SDSS).
@@
Recently, the potential of IA measurements as cosmological probes has attracted growing attention. In particular, the use of ellipticity correlations in three dimensions has been shown to enable constraints on redshift-space distortions (RSDs) and baryon acoustic oscillations (BAOs), offering both dynamical and geometric information (Chisari & Dvorkin, 2013; Okumura et al., 2019; Taruya & Okumura, 2020; Okumura & Taruya, 2022; Chuang et al., 2022; Lee et al., 2023; Saga et al., 2023; Shim et al., 2025). However, observational constraints at large scales continue to rely almost exclusively on GI correlations (Okumura & Taruya, 2023; Kurita & Takada, 2023; Xu et al., 2023b), due to the difficulty of extracting clean II signals from data.
@@
Despite the challenges, measuring shape autocorrelations remains essential because they carry unique information inaccessible to GI cross correlations. Particularly, shape autocorrelations play a critical role in probing new physics, such as the stochastic gravitational wave (GW) background, through their imprint on galaxy orientations (Schmidt et al., 2014; Biagetti & Orlando, 2020; Akitsu et al., 2023a; Okumura & Sasaki, 2024; Philcox et al., 2024; Saga et al., 2024). Recent work showed that IA signals projected along the line of sight are more naturally expanded using the associated Legendre basis rather than the standard Legendre one, due to the geometric anisotropy of projected shapes (Kurita & Takada, 2022; Okumura et al., 2024; Singh et al., 2024; Inoue et al., 2025). This formalism improves the separation of physical signals from noise and enhances interpretability at large scales.
[…]
@@ 2 SDSS and DESI Galaxy Samples
where $q$ is the minor-to-major-axis ratio ($0\leq q\leq 1$) and $\beta$ is the position angle of the ellipticity. To improve the shape measurement quality for the LOWZ and CMASS galaxies, we follow the method of Xu et al. (2023b) and crossmatch these samples with the DESI Legacy Imaging Survey DR9 data (DESI Collaboration et al., 2016; Dey et al., 2019). The DESI imaging covers the entire BOSS footprint and provides imaging data 2 to 3 magnitudes deeper than the SDSS photometry used in target selection, sufficient to robustly determine galaxy orientations. For the LRG sample, we use the original shape measurements defined by the $25~{\rm mag}~{\rm arcsec}^{-2}$ isophote in the $r$ band, which has been shown to be effective for IA analysis. Accordingly, we adopt the same LRG dataset as in Okumura & Taruya (2023).
The total numbers of the LRG, LOWZ, and CMASS used in this study are 105,334, 347,262, and 689,721, respectively. As in our previous IA analyses, we set the axis ratio to $q=0$ in Equation (5) (Okumura & Jing, 2009; Okumura et al., 2009, 2019, 2020). Since our primary interest is in the statistical accuracy of the shape correlations, rather than the amplitude of IA, this simplification does not affect our conclusions.
Figure 1: Multipole moments of the intrinsic ellipticity autocorrelation functions expanded in the standard Legendre basis, measured from SDSS LRG (left), LOWZ (middle), and CMASS (right) galaxy samples. The upper set shows the $\mathrm{II}(+)$ correlation multipoles, $\xi_{+,\ell}$, while the lower set shows the $\mathrm{II}(-)$ multipoles, $\xi_{-,\ell}$, for $\ell=0$ (monopole; circles) and $\ell=2$ (quadrupole; squares). The correlation amplitude for the LRG sample is multiplied by $1/4$. The LRG results are reproduced from Okumura & Taruya (2023), while the LOWZ and CMASS measurements incorporate high-quality DESI imaging data, significantly improving the signal. The solid curves are the best-fit nonlinear model fitted using the data points enclosed by the vertical lines. The dotted curves are the linear predictions.
@@ 3 Measurement of Correlation Functions
$\xi_{\pm}({\boldsymbol{r}})=\frac{S_{+}S_{+}\pm S_{\times}S_{\times}}{\mathrm{RR}}\ ,$ (7)
where $\mathrm{RR}$ is the normalized count of random-random pairs, and $S_{+}S_{+}=\sum_{i\neq j|{\boldsymbol{r}}}{\gamma_{+}(j|i)\gamma_{+}(i|j)}$, with $\gamma_{+}(j|i)$ being the ellipticity of galaxy $j$ measured relative to the direction to galaxy $i$, and $S_{\times}S_{\times}$ is defined similarly.
Since the IA correlations are anisotropic functions, we bin galaxy pairs in both radial separation $r=|\boldsymbol{r}|$ and the cosine of the angle to the line of sight, $\mu_{\boldsymbol{r}}$. We then compute the multipole moments of the correlation functions using the Legendre expansion,
@@
[…]
@@
where $\Theta_{L}^{m}(\mu)$ is the normalized associated Legendre function related to the unnormalized one by $\Theta_{L}^{m}(\mu)=\sqrt{\frac{2L+1}{2}\frac{(L-m)!}{(L+m)!}}\mathcal{L}_{L}^{m}(\mu_{\boldsymbol{k}})$ and $L\geq m=4$. We use a tilde to distinguish these coefficients from those expanded in the standard Legendre basis. The summation in Eqs. (8) and (9) is performed over $1\leq\mu_{\boldsymbol{r}}\leq 1$.
Figure 2: Upper panels: similarly to Fig. 1, but $L=4$ moments of $\mathrm{II}(-)$ correlation functions of SDSS galaxies expanded by the associated Legendre basis, $\widetilde{\xi}_{-,L}$. The solid curves are the best-fit nonlinear alignment and RSD models fitted using the data points enclosed by the vertical lines. Lower panels: same as upper panels, but the vertical axes are multiplied by $r^{2}$ so that they can be directly compared to the statistics shown in Fig. 1.
@@
In this analysis, we focus only on the two lowest multipoles: the monopole ($\ell=0$) and quadrupole ($\ell=2$), as the hexadecapole ($\ell=4$) is too noisy. Even for the lower multipoles, it is nontrivial to argue up to which scale the correlation is detected. The monopoles of $\mathrm{II}(+)$ appear nonzero for all three samples up to several tens of $h^{-1}\,{\rm Mpc}$, while all the quadrupoles remain statistically consistent with zero. For the $\mathrm{II}(-)$ correlation, the monopoles are better measured than those of $\mathrm{II}(+)$. The detection for LRG up to $\sim 40\,h^{-1}\,{\rm Mpc}$ corresponds to the earlier finding of Okumura et al. (2009). Notably, the LOWZ and CMASS results are significantly improved compared to those in Okumura & Taruya (2023), due to the enhanced imaging depth and quality provided by the DESI data. The quadrupole of $\mathrm{II}(-)$ for CMASS appears nonzero over a relatively large range of scales. However, because the signal is distributed over multiple multipoles and the overall measurements remain noisy, these results do not allow a definitive detection based solely on the standard Legendre expansion.
@@ 4.1 Nonlinear Alignment and RSD Model
Table 1: Constraints on the Amplitude of IA and Reduced $\chi^{2}$ Values.
@@
Statistics  dof $b_{K}\sigma_{8}$ $\chi_{\min}^{2}/{\rm dof}$ $b_{K}\sigma_{8}$ $\chi_{\min}^{2}/{\rm dof}$ $b_{K}\sigma_{8}$ $\chi_{\min}^{2}/{\rm dof}$
@@
where $\delta_{m}$ is the nonlinear matter density field and $b_{K}$ is the redshift-dependent coefficient of the IA, which we refer to as the shape bias. [^1] The shape bias parameter is related to $\widetilde{C}_{1}$ in Okumura & Taruya (2020) by $b_{K}=-\widetilde{C}_{1}$. Furthermore, the redshift-space shape field is multiplied by the damping function due to the Finger-of-God (FoG) effect. We then define EE-/$B$-modes, $\gamma_{(E,B)}$, which are the rotation-invariant decomposition of the ellipticity field (Crittenden et al., 2002),
@@
where $k=|\boldsymbol{k}|$, $\mu_{\boldsymbol{k}}$ is the direction cosine between the line of sight and $\boldsymbol{k}$, and $P_{\delta\delta}$ is the nonlinear matter power spectrum computed by the revised Halofit (Takahashi et al., 2012). Since the FoG effect enters IA power spectra in the same way as the conventional galaxy power spectrum, we adopt a simple Gaussian function to characterize the FoG damping, $D_{\rm FoG}(k\mu_{\boldsymbol{k}}\sigma_{v})=\exp{\left(-k^{2}\mu_{\boldsymbol{k}}^{2}\sigma_{v}^{2}/2\right)}$, with $\sigma_{v}$ the nonlinear velocity dispersion parameter. Since $P_{\delta\delta}$ is proportional to the square of the normalization parameter of the density fluctuation, $\sigma_{8}^{2}(z)$, free parameters for this model are ${\boldsymbol{\theta}}=(b_{K}\sigma_{8},\sigma_{v})$.
To take into account the line-of-sight anisotropy in Equation (16), we use the multipole expansion in terms of the standard or associated Legendre polynomials (Kurita & Takada, 2022; Okumura et al., 2024),
@@
$\xi_{+,\ell}(r)=i^{\ell}\int\frac{k^{2}dk}{2\pi^{2}}j_{\ell}(kr)P_{EE,\ell}(k)~,$ (19)
$\widetilde{\xi}_{-,L}(r)=i^{L}\int\frac{k^{2}dk}{2\pi^{2}}j_{L}(kr)\widetilde{P}_{EE,L}(k)~,$ (20)
where $j_{\ell}$ is the spherical Bessel function. Converting the $\mathrm{II}(-)$ correlation function from the associated to the standard Legendre bases can be performed via the linear transform,
@@
where the coefficients $A_{\ell L}$ are constants and provided in Appendix A. In the linear-theory limit, $\sigma_{v}\to 0$ and thus $D_{\rm FoG}=1$. Hence, Equation (16) converges to the linear-theory formulas derived in Okumura & Taruya (2020) (see also Appendix C of Okumura et al., 2024).